OpenStudy (anonymous):

Given a cube with a volume of 36 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? 3cm 6cm-NO 9cm 12cm -No

jimthompson5910 (jim_thompson5910):

Volume of Pyramid = (1/3)*(Volume of Cube)

OpenStudy (anonymous):

2) Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi. C = 5π; A = 20π C = 10π; A = 20π C = 10π; A = 25π- i think C = 5π; A = 25π

jimthompson5910 (jim_thompson5910):

C = 10π; A = 25π is correct

OpenStudy (anonymous):

3) The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 360 mm3, what is the volume of the stack of pennies on the right in cubic millimeters? Use 3.14 for pi.

OpenStudy (anonymous):

@jim_thompson5910 so would number one be 9cm

jimthompson5910 (jim_thompson5910):

Volume of Pyramid = (1/3)*(Volume of Cube) Volume of Pyramid = (1/3)*(36) Volume of Pyramid = 12 cubic cm

jimthompson5910 (jim_thompson5910):

So there has to be a typo somewhere

OpenStudy (anonymous):

i tried 12 and 6 both where wrong pretest

OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 10 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest tenth

jimthompson5910 (jim_thompson5910):

then the next closest thing is 9 but I still think there's a typo

OpenStudy (anonymous):

what about number three

OpenStudy (anonymous):

and the last one

OpenStudy (mathmale):

Jim T.: I'd appreciate hearing the reasoning behind "Volume of Pyramid = (1/3)*(Volume of Cube)." Vol of a pyramid =(1/3) (area of base) (height) Vol of a cube:= (area of base) (height). Could you cite some reference that testifies to and explains this?

jimthompson5910 (jim_thompson5910):

if you look at this page http://www.mathopenref.com/pyramidvolume.html it says V = (1/3)*b*h V = volume of pyramid b = area of base h = height but b*h is the volume of the prism with the same base and height as the pyramid

OpenStudy (anonymous):

please i only need number 3 and the last one

OpenStudy (mathmale):

3) The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 360 mm3, what is the volume of the stack of pennies on the right in cubic millimeters? Use 3.14 for pi. Still there, Britt? And, if so, is this the problem you want help with?

OpenStudy (mathmale):

Unless I'm sorely mistaken, the answer is unbelievably simple, and much of the info given in this problem is irrelevant and unneeded. Think about this problem that way and see what you come up with.

OpenStudy (anonymous):

yes i am here

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

148.365

OpenStudy (mathmale):

Britt: What is the volume of the stack of pennies on the left?

OpenStudy (anonymous):

that what i got i did it wrong i don't understand

OpenStudy (mathmale):

britt5 typed: 3) The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 360 mm3

OpenStudy (mathmale):

then, again, what is the volume of the left stack?

OpenStudy (mathmale):

Britt: the problem statement, as you typed it, tells us that the volume of the left stack of pennies is 360 mm^3 (360 cubic millimeters).

OpenStudy (mathmale):

Britt, have you any reason to think that the stack of pennies on the right is different in volume from that on the left?

OpenStudy (anonymous):

just the picture

OpenStudy (mathmale):

But you could straighten up that stack of pennies, and then that stack on the right would be identical to the stack on the left, wouldn't it? So try again to define/calculate/find the volume of the stack on the right.

OpenStudy (anonymous):

thank @mathmale

OpenStudy (mathmale):

Britt, the point I'm trying to make is that the 2 stacks of pennies have precisely the same volume. so the answer , in cubic mm, is simply 360 mm^3, same as for the left stack. OK?

OpenStudy (anonymous):

Given a cube with a volume of 36 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? is this questions answer 9????